OpenStudy (anonymous):

write an equation of the line that passes through the point an has the given slope. write the equation in slope intercept form (-3,-5) m=-3/5

OpenStudy (anonymous):

y=mx+b -5=-3/5(-3)+b -5=1.8+b -6.8=b y=-3/5x -6.8